# From Trace to Line: LLM Agent for Real-World OSS Vulnerability Localization

### v1

@@ 4.4 T2L-ARVO Benchmark
The ARVO dataset contains over 4,900 reproducible vulnerabilities across 250+ C/C++ projects but lacks the structure needed to evaluate agentic for vulnerability localization. We introduce T2L-ARVO, a 50-case benchmark with comprehensive crash-type coverage and graded difficulty for LLM agent evaluation. ARVO’s human-oriented, reproducible builds require adaptation for agentic assessment; we therefore apply a dual validation layer—manual expert checks plus LLM-assisted verification—to ensure selected cases are both faithfully reproducible and appropriately challenging for automated agents. This yields a benchmark with realistic, graded difficulty and broad crash-type coverage, enabling rigorous, end-to-end evaluation of planner–executor systems for trace-to-line vulnerability localization—well beyond single-project or single-crash-type studies.
@@
ARVO Analysis We analyzed 4,9934{,}993 ARVO instances and grouped them by underlying failure mechanism: Buffer Overflows $49.9\%$ (n=2,490)(n=2{,}490), Uninitialized Access & Unknown States $35.4\%$ (n=1,768)(n=1{,}768), Memory Lifecycle Errors $11.5\%$ $(n=573)$, Type Safety & Parameter Validation $2.9\%$ $(n=147)$, and System & Runtime Errors $0.3\%$ $(n=15)$.
Each family subsumes concrete subtypes (e.g., heap-buffer-overflow, use-of-uninitialized-value, heap-use-after-free, bad-cast). T2L-ARVO deliberately mirrors this distribution to avoid bias toward any single failure mode.
Verification Process We combine automated screening with expert review to ensure both realism and balance. Quantitatively, we score candidates using diff-based structural metrics (files changed, architectural spread, directory depth) and semantic factors (cross-module coupling, interface changes, concurrency touchpoints). We then apply dual validation: (i) manual expert assessment to confirm reproducibility and representativeness, and (ii) LLM-assisted checks to gauge agent-facing difficulty. Known patch locations are indexed to establish clear localization baselines and to support precise, line-level scoring.
T2L-ARVO Composition. The final benchmark comprises 50 vulnerabilities, evenly sampled across five crash families (10 each) for broad yet controlled difficulty. Each family includes representative subtypes (e.g., heap-buffer-overflow, use-of-uninitialized-value, heap-use-after-free, bad-cast), covering single-file defects and cross-module interactions to prevent overflow bias and exercise diverse failure modes observed in real repositories.
@@ 5 Experiment Setup
Implementation. We build T2L-Agent from scratch without LangChain, DSPy, and LlamaIndex to keep the core lightweight, retain fine-grained control over reasoning and tools, and maximize extensibility. The Agentic Trace Analyzer compiles targets with ASAN and collects crashes, stack traces, and allocation metadata to yield actionable traces for narrowing. Following ARVO’s layout, we maintain registry-backed, per-project environments and provide both vulnerable and patched revisions in containers. Our harness runs dockerized T2L-Agent’s that interface with T2L-ARVO via the Docker SDK for Python, orchestrating build–run–reproduce cycles from within the agent loop—ensuring consistent conditions, deterministic reproduction, and auditable measurement.
@@ 6.1 Baseline Benchmarking
Table 1: Localization and Detection Rate Performance Across Different Models.
@@
Taken together, equal budgets surface clear profiles. GPT-5 (low-think) converts evidence into the strongest line-level localization, while GPT-4.1 extracts slightly more coarse-grained signal at the chunk level. GPT-4o-mini’s mix—competitive detection (44.3%) but weak localization (22.6%)—suggests higher recall with looser ranking that does not always translate to precise line hits. Open-source models lag on both metrics under the same constraints, indicating gaps in code understanding and tool use rather than simple parameter tuning. Overall, improvements track the availability of concrete runtime evidence, and structured, tool-grounded reasoning appears more impactful than generation settings for end-to-end vulnerability localization.
@@ 6.2 Discussion 1: Feature-wise Evaluation
Agentic Trace Analyzer. This table  2 demonstrates the critical effectiveness of our proposed Agentic Trace Analyzer (ATA) through ablation experiments. Without ATA, both GPT-5 and Claude 4 Sonnet achieve 0.0% detection and localization rates across all vulnerability families. This complete performance breakdown validates that our ATA component successfully bridges the gap between crash symptoms and vulnerability locations, addresses the fundamental challenge of vulnerability localization in complex codebases.
@@
Table 2: Localization and Detection Rate Performance with Refinement and Divergence Tracing Across Different Models.
@@
Divergence Tracing. Tab.  2 shows divergence tracing delivers the strongest overall gains over baseline. All models improve on both metrics: GPT-5 is up 13.7% in detection and 10.3 in localization; GPT-4.1 gains 4.0% and 11.4% respectively. Qwen3 Next see the largest jumps with adding 13.8% in detection and 48.9% in localization, while Qwen3 235B adds 16.8% and 32.9%. These consistent lifts across architectures highlight divergence tracing as a core algorithmic upgrade for vulnerability localization.
@@ 6.3 Discussion 2: Parameter Tuning
Table 3: Localization and Detection Rate Performance for Temperature Tuning and Chain of Thought Across Different Models.
@@ 8 Conclusion
To ensure reproducibility, we will publicly release the full T2L-Agent framework upon publish of this paper, including all module codebase, evaluation scripts, and benchmark data used in this paper. Our implementation does not rely on any proprietary components. Detailed descriptions of our methodology are provided in the main text (Sections 4.1, 4.4) and supported by step-by-step examples in the appendix. All experimental configurations, including model versions, prompting strategies, and budget constraints, will be documented and made available upon publication to enable full replication of our results.

### v2

@@ 4.3 T2L-ARVO Benchmark
The original ARVO dataset contains over 4,993 reproducible vulnerabilities across 250+ C/C++ projects, but its human-oriented design and build-centric structure do not directly support evaluating agentic vulnerability-localization systems. To bridge this gap, we developed T2L-ARVO, a 50-case benchmark derived from ARVO for assessing LLM agents in evidence-guided vulnerability localization, rather than human bug reproduction.T2L-ARVO maintains a moderate, calibrated difficulty while keeping the original crash-type diversity. Its construction employs a dual validation pipeline combining manual expert review with LLM-assisted checks to ensure each case is reproducible and appropriately challenging for automated agents. This yields a realistic, balanced benchmark for rigorous end-to-end evaluation of LLM-based vulnerability localization. In T2L-ARVO, each of the cases has multiple ground-truth bug lines. The average accuracy overall is denoted as $\textit{Accuracy}=\textit{correct localized vulnerable lines}/\textit{total vulnerable lines}$.
T2L-ARVO Composition We analyzed 4,9934{,}993 ARVO instances and grouped them by underlying failure mechanism: Buffer Overflows $49.9\%$ (n=2,490)(n=2{,}490), Uninitialized Access & Unknown States $35.4\%$ (n=1,768)(n=1{,}768), Memory Lifecycle Errors $11.5\%$ $(n=573)$, Type Safety & Parameter Validation $2.9\%$ $(n=147)$, and System & Runtime Errors $0.3\%$ $(n=15)$. Each family subsumes concrete subtypes (e.g., heap-buffer-overflow, use-of-uninitialized-value, heap-use-after-free, bad-cast). T2L-ARVO deliberately mirrors this distribution to avoid bias toward any single failure mode.
The final benchmark comprises 50 vulnerabilities samples, evenly sampled across five crash families (10 each) for broad yet controlled difficulty. Each family includes representative subtypes (e.g., heap-buffer-overflow, use-of-uninitialized-value, heap-use-after-free, bad-cast), covering single-file defects and cross-module interactions to prevent overflow bias and exercise diverse failure modes observed in real repositories.
@@
Verification Process Our verification follows two main considerations. First, ARVO crash families contain multiple subtypes. To reduce redundancy and improve representativeness, we select only a few typical subtypes from each family (see A.1), preserving diversity without overloading the benchmark with similar vulnerabilities. Second, experts evaluate each sample’s complexity to keep T2L-ARVO at a moderate, calibrated difficulty. We score candidates using diff-based structural metrics (e.g., files changed, architectural spread, directory depth) and semantic factors (e.g., cross-module coupling, interface changes), ensuring that T2L-ARVO remains balanced in both coverage and difficulty.
@@ 5 Experiment Setup
Implementation. We build T2L-Agent from scratch without LangChain, DSPy, and LlamaIndex to keep the core lightweight, retain fine-grained control over reasoning and tools, and maximize extensibility. The Agentic Trace Analyzer compiles targets with ASAN and collects crashes, stack traces, and allocation metadata to yield actionable traces for narrowing. Following ARVO’s layout, we maintain registry-backed, per-project environments and provide both vulnerable and patched revisions in containers. Our harness runs dockerized T2L-Agent’s that interface with T2L-ARVO via the Docker SDK for Python, orchestrating build–run–reproduce cycles from within the agent loop, ensuring consistent conditions, deterministic reproduction, and auditable measurement.
@@ 6.1 Baseline Benchmarking
Table 1: Base T2L-Agent Localization and Detection Rate Performance Across Different Models.
@@
Taken together, equal budgets surface clear profiles. GPT-5 (low-think) converts evidence into the strongest line-level localization, while GPT-4.1 extracts slightly more coarse-grained signal at the chunk level. GPT-4o-mini’s mix—competitive detection (44.3%) but weak localization (22.6%), suggests higher recall with looser ranking that does not always translate to precise line hits. Open-source models lag on both metrics under the same constraints, indicating gaps in code understanding and tool use rather than simple parameter tuning. Overall, improvements track the availability of concrete runtime evidence, and structured, tool-grounded reasoning appears more impactful than generation settings for end-to-end vulnerability localization.
@@ 6.2 Discussion 1: Feature-wise Evaluation
Agentic Trace Analyzer. This table  2 demonstrates the critical effectiveness of our proposed Agentic Trace Analyzer (ATA) through ablation experiments. Without ATA, GPT-5 and Claude 4 Sonnet show 0.0% detection and localization across all families. This performance breakdown validates that our ATA component successfully bridges the gap between crash symptoms and vulnerability locations, addresses the fundamental challenge of vulnerability localization in complex codebases.
@@
Table 2: Localization and Detection Rate Performance of T2L-Agent with Refinement and Divergence Tracing Across Different Models.
@@
Divergence Tracing. Tab. 2 shows divergence tracing yields the strongest gains. All models improve on both metrics: GPT-5 is up 13.7% in detection and 10.3 in localization; GPT-4.1 gains 4.0% and 11.4% respectively. Qwen3 Next see the largest jumps with adding 13.8% in detection and 48.9% in localization, while Qwen3 235B adds 16.8% and 32.9%. These consistent lifts across architectures highlight divergence tracing as a core algorithmic upgrade for vulnerability localization.
@@ 6.3 Discussion 2: Parameter Tuning
Table 3: Localization and Detection Rate Performance of T2L-Agent for Temperature Tuning and Chain of Thought Across Different Models.
@@ 8 Conclusion
To ensure reproducibility, we will publicly release the full T2L-Agent framework upon publish of this paper, including all module codebase, evaluation scripts, and benchmark data used in this paper. Our implementation does not rely on any proprietary components. Detailed descriptions of our methodology are provided in the main text (Sections 4.1, 4.3) and supported by step-by-step examples in the appendix. All experimental configurations, including model versions, prompting strategies, and budget constraints, will be documented and made available upon publication to enable full replication of our results.